OpenStudy (anonymous):

A movie theatre has a capacity of 49 seats. When there are d spectators in the theatre, there are z+13 empty seats. With t spectators, there are z+ 19 empty seats. What is the values of c?

OpenStudy (anonymous):

what is c here?

OpenStudy (anonymous):

A movie theatre has a capacity of 49 seats. When there are d spectators in the theatre, there are z+13 empty seats. With c spectators, there are z+ 19 empty seats. What is the values of c?

OpenStudy (anonymous):

its c spectators not t sorry

OpenStudy (anonymous):

c = d-6 or 30-z you want any specific number?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

49= (z) + (z+17) 49= (c) + (z +11)

OpenStudy (anonymous):

Actually, it can be 49 = c + (z + 19) or 49 = d + (z + 13)

OpenStudy (anonymous):

can you solve one so i can see if i'm right